OpenStudy (anonymous):

how can you multiply (7/3)(7/3) / (7/3)'square? is the answer 14/3 or 49/3? reply back asap! thnx :)

OpenStudy (anonymous):

49/9

OpenStudy (anonymous):

thanks :D

jhonyy9 (jhonyy9):

no this will be (49/9)/(49/9)=(49/9)(9/49)=1